Matt's Eat 'N Run has two categories of direct labor: unskilled, which costs $10 per hour, and skilled, which costs $20 per hour. Management has established standards per "equivalent friendly meal," which has been defined as a typical meal consisting of a hamburger, a drink, and french  fries. Standards have been set as follows.

 




Skilled labor2minutes per equivalent meal
Unskilled labor6minutes per equivalent meal

 

For the year, Matt’s sold 180,000 equivalent friendly meals and incurred the following labor costs.

 





Skilled labor6,000hours
$125,000
Unskilled labor15,000hours

240,000

 

Required:
a.
 Compute labor price and efficiency variances.
b. Compute labor mix and yield variances.
Answer #1

A) 

Skilled:
(AP × AQ) = $125,000
(SP × AQ) = $20 × 6,000 = $120,000
Labor price variance: $120,000 – $125,000 = $5,000 U

Labor efficiency variance: $120,000 – $120,000 = $0 None


Unskilled:
(AP × AQ) = $240,000
(SP × AQ) = $10 × 15,000 = $150,000
Labor price variance: $150,000 – $240,000 = $90,000 U

Labor efficiency variance: $180,000 – $150,000 = $30,000 F


Total
(AP × AQ) = $365,000
(SP × AQ) = $270,000
Labor price variance: $270,000 – $365,000 = $95,000 U

Labor efficiency variance: $300,000 – $270,000 = $30,000 F

                                       

                                         Skilled                          Unskilled                    Total

Labor price variance       5,000 U                         90,000U                95,000U  Labor efficiency variance                                       30,000F                30,000F


B) 

Skilled:
(SP × ASQ) = $20 × (0.25 × 21,000) = $20 × 5,250 = $105,000
(SP × SQ) = $20 × (2/60 × 180,000) = $20 × 6,000 = $120,000
Labor mix variance: $105,000 – $120,000 = $15,000 U
Labor yield variance: = $120,000 – $105,000 = $15,000 F

Unskilled:
(SP × ASQ) = $10 × (0.75 × 21,000) = $10 × 15,750 = $157,500
(SP × SQ) = $10 × (6/60 × 180,000) = $10 × 18,000  = $180,000
Labor mix variance: $157,500 – $150,000 = $7,500 F
Labor yield variance: = $180,000 – $157,500 = $22,500 F

Total:
(SP × ASQ) = $262,500
(SP × SQ) = $300,000
Labor mix variance: $262,500 – $270,000 = $7,500 U

Labor yield variance: $300,000 – $262,500 = $37,500 F


SkilledUnskilledTotal
Labor mix variance$15,000$7,500$7,500
Labor yield variance$15,000$22,500$37,500